OpenStudy (anonymous):

5^(2x-3)=25^(3x+1) solve for x

OpenStudy (hoblos):

first you must know that ln(a^x) = xln(a) so if we introduce the ln here for both sides ln{5^(2x-3)}=ln{25^(3x+1)} (2x-3)ln5 = (3x+1)ln25 (2x-3)ln5 = (3x+1)ln(5^2) (2x-3)ln5 = 2(3x+1)ln5 divide by ln5 2x-3 = 2(3x+1) 2x-3 = 6x +2 -4x = 5 x = -5/4

OpenStudy (anonymous):

Wow, thanks, over teacher didn't show us this rule so I didn't know how to solve.

OpenStudy (anonymous):

what happens when I have ln16 and ln8? how do i cancel those?

OpenStudy (hoblos):

ln16 = ln(2^4) = 4ln2 ln8 = ln(2^3) = 3ln2

OpenStudy (anonymous):

ohh I see. Thanks!

OpenStudy (hoblos):

any time :)